OpenStudy (anonymous):

help pweassssssssse In 2010, the 2009 populations and projected population rates were released for Russia and Ethiopia. The population of Russia in 2009 was 140.9 million; the population of Ethiopia was 82.8 million. Russia is decreasing in population by approximately 0.4% annually; the population of Ethiopia is growing at 2.6%. Using logs and letting x be the number of years since 2009, answer the following questions.

OpenStudy (anonymous):

a. How long, at the above rates, will it take for the population of Ethiopia to double? b. How long, at the above rates, will it take for the population of Russia to reach 125 million?

OpenStudy (anonymous):

we use the excel fo problems like this btw

OpenStudy (anonymous):

for*

OpenStudy (anonymous):

This is similar to our other question. For Ethiopia, you don't even need a number for the population though. Let initial population be P and eventual population be 2P.

OpenStudy (anonymous):

I have to take off... good luck... glad to see you're in good hands with @tcarroll010

OpenStudy (anonymous):

P(1.026)^x = 2P. Or, more simply, (1.026)^x = 2. Can you do logs now the way I showed you?

OpenStudy (anonymous):

b is around 2038

OpenStudy (anonymous):

what do u insert in replacement of x i dont understand that part

OpenStudy (anonymous):

I didn't get to b yet, I'm still on a. You young folk are so fast! x(log(1.026)) = log(2).

OpenStudy (anonymous):

x is the passage of years in number of years.

OpenStudy (anonymous):

actually around 2039 would be more accurate for b but i have no clue how to use logs

OpenStudy (anonymous):

a is definietly a problem for me

OpenStudy (anonymous):

27.005 years for part a. Do you agree?

OpenStudy (anonymous):

can u go thru the steps with me please? i hate to admit it but im a bit slow with the logs

OpenStudy (anonymous):

absolutely no problem. A lot of math is letting things sink in. You're not expected to be an expert right off. But to learn, you will have to get the concepts down after reviewing the steps and answers which is what you should do when anyone helps you. Try to do the whole problem again yourself if you got lots of help.

OpenStudy (anonymous):

For Ethiopia, we are being asked how long does it take for the population to double, so it doesn't really matter what the initial population is. The concept will work for whatever the initial population is.

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

So, initial pop is P and eventual pop is 2P. This will simplify things. You could work with the actual numbers but that muddies up what is being asked. Clearer this way, you'll see.

OpenStudy (anonymous):

82.8(1+.026) is this used at all

OpenStudy (anonymous):

That's a good and commendable attempt and will lead to the right answer, but better and more mathematically elegant to say P(1.026) after the first year and P(1.026)^2 after the second and P(1.026)^x after the xth year.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

how would we apply this to the logs

OpenStudy (anonymous):

You are then looking for P(1.026)^x >= 2P

OpenStudy (anonymous):

2P because we're looking fo the doubled amount?

OpenStudy (anonymous):

Here's where the elegance and clarity come into play. With P and 2P you are driving home the concept of doubling. Also >= because you want to "get to" 2P and the > in >= emphasizes that you are looking for a minimum value for x. Yes, 2P for doubled amount.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

And the beauty of P(1.026)^x >= 2P shows 2 things. 1) emphasizing doubling, and 2) that the initial value of P doesn't matter because of the obvious cancellation. So, you get (1.026)^x >= 2. Simpler is better.

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

Since x is in the exponent, logs are needed to solve. So, taking logs of both sides, we get x(log(1.026)) = log(2).

OpenStudy (anonymous):

Sorry, x(log(1.026)) >= log(2).

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

im writing all this down

OpenStudy (anonymous):

Easy solving logs and simple division gives 27.005 rounded.

OpenStudy (anonymous):

Take your time. Of course, you always have access to this question since you own it and can get to it.

OpenStudy (anonymous):

is there division involved?

OpenStudy (anonymous):

That last equation, x(log(1.026)) >= log(2), becomes x >= log(2) / log(1.026). That's the division.

OpenStudy (anonymous):

With a numerical answer of 27.005, you can how ridiculous it would be to multiply (1.026) by itself 27+ times to get to 2. But, instead of multiplying 27+ times, you could just use your calculator to try (1.026)^27.005 to see how this works.

OpenStudy (anonymous):

2

OpenStudy (anonymous):

Yes, and nice that logs exists that allow us to work backward.

OpenStudy (anonymous):

so it would only take 2 years to double their population?..

OpenStudy (anonymous):

wait that doesnt sound likely

OpenStudy (anonymous):

No, 27.005 to double. 2 is the "twice"

OpenStudy (anonymous):

ohhhhhh oh ok i see

OpenStudy (anonymous):

P(1.026)^27.005 = 2P. x is the years and is 27.005

OpenStudy (anonymous):

yes i see now

OpenStudy (anonymous):

Don't worry, you'll get it. It does take some time. Logs and exponents is where I see a lot of people get confused at the start. But just stick with it.

OpenStudy (anonymous):

so did i do b correctly?

OpenStudy (anonymous):

You appeared to get an answer of about 29 years for b? Is that right? I didn't do b yet, but I could do it and check.

OpenStudy (anonymous):

yes can u check for me i got the year 2039

OpenStudy (anonymous):

am i supposed to use logs for this cuz i didnt i used a simpler method

OpenStudy (anonymous):

140.9(0.996)^x <= 125 is the formula

OpenStudy (anonymous):

Well, whatever is easiest for you. I use logs because that is what is easiest for me.

OpenStudy (anonymous):

yes thats what i used lol i guess i was right afterall then

OpenStudy (anonymous):

logs r so difficult for me i need alot more practice

OpenStudy (anonymous):

x(log(0.996) <= log(0.88715)

OpenStudy (anonymous):

then u divide that?

OpenStudy (anonymous):

x is 29.874

OpenStudy (anonymous):

Yes. How did you do this without logs?

OpenStudy (anonymous):

wait ... i think i did it wrong then

OpenStudy (anonymous):

i did 140.9(1-.004)

OpenStudy (anonymous):

Well, you're very close and since you approximated, I'd say you got an ok answer with year 2039 because that is about 30 years from 2009. (1 - 0.004) is the same as my 0.996 so you are ok with that.

OpenStudy (anonymous):

Didn't you use (1 - 0.004)^x somewhere?

OpenStudy (anonymous):

x(log(0.996) <= log(0.88715)

OpenStudy (anonymous):

do u divide that

OpenStudy (anonymous):

Yes, like the part a. you divide out the log expression on the left to isolate x.

OpenStudy (anonymous):

so log0.996/log0.88715= 29.874?

OpenStudy (anonymous):

Just one change in my equation. I should have put >= not <=. As for your question, switch the numerator and denominator.

OpenStudy (anonymous):

so log0.88715/log0.996= 29.874?

OpenStudy (anonymous):

So, with that, you're done!

OpenStudy (anonymous):

hmmm still need practice but thank u so much

OpenStudy (anonymous):

You'll get it. Just practice and review.

OpenStudy (anonymous):

kk thank u

OpenStudy (anonymous):

you're welcome